Suppose that the probability of a baseball player getting a hit in an at-bat is 0.2804. If the player has 33 at-bats during a week, what's the probability that he gets no more than 9 hits?

Question 6 options: 1) 0.0768 2) 0.1536 3) 0.5493 4) 0.4507 5) 0.3957
Answer #1

X ~ Binomial (n,p)

P(X) = nCx px (1-p)n-x

P( X <= 9) = P( X = 0) + P( X = 1) + P( X = 2) + P( X = 3) + P( X = 4) + P( X = 5) + P( X = 6) + P( X = 7) +

P( X = 8) + P( X = 9)

=  33C0 0.28040 0.719633 + 33C1 0.28041 0.719632 + 33C2 0.28042 0.719631 + 33C3 0.28043 0.719630 +

33C4 0.28044 0.719629 + 33C5 0.28045 0.719628 + 33C6 0.28046 0.719627 + 33C7 0.28047 0.719626 +

33C8 0.28048 0.719625 + 33C9 0.28049 0.719624  

= 0.5493